UT 2012-07-30
=============

CHARA/CLIMB

B1   B2   B3   B4   B5   B6
S2   E2   W1
     E1   W2

S1: POP 5
S2: POP 5
E1: POP 1
E2: POP 1
W1: POP 4
W2: POP 5

seeing 10-15 cm


*** S2-E2-W1

UT 03:40 - hd144585 (fringe finder)
   	   doing NIRO alignment
	   found fringes

UT 04:15 - hd145518
   	   doing NIRO alignment
   	   recording LONG, 500 Hz, NONDEST

UT 04:25 - AS205A
   	   recording LONG, 500 Hz, NONDEST
	   in the beginning only fringe 12 there, later all 3 -> good quality
   	   saturated sometimes slightly (USE TRUNCATE FUNCTION DURING DATA REDUCTION TO REMOVE IT!)

UT 04:55 - hd145518
   	   recording LONG, 500 Hz, NONDEST, sometimes noisy on B23, B31
	   flux from one beam is low, so we redo NIRO alignment
   	   recording LONG, 750 Hz, NONDEST, much better!

UT 05:30 - AS205A
   	   recording LONG, 750 Hz, NONDEST, good data on all three beams

UT 05:45 - hd148198
   	   started recording (recording LONG, 750 Hz, NONDEST), but then OPLE computer dies
	   tried to fix it, but it's a serious problem with the motherboard.
	   The beep signal from the motherboard (and some popped capacitors) indicates that 
	   board is damaged, which we cannot repair during the night.